# A Practical Introduction to Kernel Discrepancies: MMD, HSIC & KSD

### v1

@@ Abstract
This paper corresponds to the introduction of my PhD thesis (Schrab, 2025a, Chapter 2) and is presented as a standalone article to introduce the reader to kernel discrepancies estimators. First, in Section 1, we define kernels, Reproducing Kernel Hilbert Spaces, mean embeddings and cross-covariance operators, and present kernel properties such as characteristicity, universality and translation invariance. Then, in Section 2, we introduce the Maximum Mean Discprecancy, the Hilbert–Schmidt Independence Criterion, and the Kernel Stein Discrepancy, as well as their estimators, and we discuss the importance of the choice of kernel for such measures. We then introduce a collection of statistics in Section 3, including the commonly-used complete statistics, as well as their incomplete counterparts which trade accuracy for computational efficiency. Finally, in Section 4, we construct adaptive estimators combining multiple statistics with various kernels, which is one method to address the problem of kernel selection.
@@ 1 Kernels
  • $[1\Longrightarrow 2]$ Every kernel is a positive definite kernel as $\sum_{i=1}^{n}\sum_{j=1}^{n}c_{i}c_{j}k(x_{i},x_{j})=\|\sum_{i=1}^{n}c_{i}\phi(x_{i})\|_{{\mathcal{H}}}\geq 0$.
  • $[2\Longrightarrow 1]$ Every positive definite kernel is guaranteed to be an inner product between features in an Hilbert space (Steinwart and Christmann, 2008, Theorem 4.16).
  • $[2\Longrightarrow 3]$ Given a positive definite kernel $k$ on ${\mathcal{X}}\times{\mathcal{X}}$, there exists a unique Hilbert space ${\mathcal{H}}_{k}$ of real-valued functions on ${\mathcal{X}}$ for which $k$ is the reproducing kernel, ${\mathcal{H}}_{k}$ is the (unique) RKHS associated to $k$ (Aronszajn, 1950, Moore-Aronsajn Theorem).
  • $[3\Longrightarrow 1]$ A reproducing kernel is a kernel with feature map $\phi(x)=k(x,\cdot)$ as $\langle k(x,\cdot),k(y,\cdot)\rangle_{{\mathcal{H}}}=k(x,y)$.
@@
  • $[3\Longrightarrow 4]$ Using the reproducing property followed by Cauchy–Schwartz inequality, we obtain $|f(x)|\leq\sqrt{k(x,x)}\|f\|_{{\mathcal{H}}}$ for all $f\in{\mathcal{H}}$ and all $x\in{\mathcal{X}}$.
  • $[4\Longrightarrow 3]$ By Riesz Representation Theorem (Rudin, 1987, Theorem 4.12), for all $x\in{\mathcal{X}}$ there exists $g_{x}\in{\mathcal{H}}$ such that $f(x)=\langle f,g_{x}\rangle_{{\mathcal{H}}}$ for all $f\in{\mathcal{H}}$. The function $k(x,y)\coloneqq g_{x}(y)$ satisfies the two properties of definition 3 and hence is a repoducing kernel.
@@
=\ = $\mathbb{E}_{P_{XY}}\!\left[\left\langle f,\Big{(}\big{(}{k^{\mathcal{X}}}(X,\cdot)-\mu_{P_{X}}\big{)}\otimes\big{(}{k^{\mathcal{Y}}}(Y,\cdot)-\mu_{P_{Y}}\big{)}\Big{)}g\right\rangle_{{\mathcal{H}}_{k^{\mathcal{X}}}}\right]$
@@ 2 Kernel discrepancies
We introduce the Maximum Mean Discrepancy (MMD) in Section 2.1, the Hilbert–Schmidt Independence Criterion (HSIC) in Section 2.2, and the Kernel Stein Discrepancy (KSD) in Section 2.3.
@@ 2.1 MMD: Maximum Mean Discrepancy
We now introduce some estimators of the MMD given some independent samples $X_{1},\dots,X_{m}\overset{\mathrm{i.i.d.}}{\sim}P$ and $Y_{1},\dots,Y_{n}\overset{\mathrm{i.i.d.}}{\sim}Q$. We let $\widehat{P}$ and $\widehat{Q}$ denote the empirical distributions (uniform distributions on the datapoints). The plug-in estimator (Gretton et al., 2012, Equations 2 and 5) for $\mathrm{MMD}^{2}_{k}(P,Q)$ is $\mathrm{MMD}^{2}_{k}(\widehat{P},\widehat{Q})$ which from Equation 6 is equal to
@@
for $x,x^{\prime},y,y^{\prime}\in\mathbb{R}^{d}$. Writing the estimator $V_{\mathrm{MMD}^{2}_{k}}$ as a two-sample V-statistic can be theoretically appealing but we stress that it can in fact be computed in quadratic time using Equation 7. The V-statistic incorporates the terms $\{k(X_{i},Y_{i}):i=1,\dots,n\}$ and, hence, is biased. We also point out that the V-statistic is always non-negative, and taking its square root yields an estimator of the (non-squared) MMD. The global sensitivity of this MMD statistic is studied in Kim and Schrab (2023, Lemma 5), which provides robustness guarantees (Schrab and Kim, 2025).
@@
An unbiased estimator of the squared MMD (Gretton et al., 2012, Lemma 6) naturally arises from Equation 6 as
@@
The MMD U-statistic expression of Equation 12 cannot be expressed as a single vector-matrix-vector product, instead it needs to be computed as
@@
$\widetilde{U}_{\mathrm{MMD}^{2}_{k}}=\frac{1}{n(n-1)}\bm{1}^{\top}\bar{K}^{ZZ}\bm{1}$ (16)
where $\bar{K}^{ZZ}$ is the matrix $K^{ZZ}$ as defined in Equation 10 but with diagonal entries set to zero. As a result of the U-statistic being unbiased, it is not always non-negative and hence cannot be used to estimate the MMD by taking its square root (as for the V-statistic).
@@
When using the translation-invariant kernel $k_{\lambda}$, Equation 3 implies [^3] We use the convention that $(\cdot)_{\lambda}$ denotes $(\cdot)_{k_{\lambda}}$ throughout the thesis.
@@ 2.2 HSIC: Hilbert–Schmidt Independence Criterion
=\ = $\big{\langle}{\mathcal{C}}_{P_{XY}},{\mathcal{C}}_{P_{XY}}\big{\rangle}_{\mathrm{HS}}$
=\ = $\big{\langle}\mathbb{E}_{P_{XY}}\!\left[\big{(}{k^{\mathcal{X}}}(X,\cdot)-\mu_{P_{X}}\big{)}\otimes\big{(}{k^{\mathcal{Y}}}(Y,\cdot)-\mu_{P_{Y}}\big{)}\right],\mathbb{E}_{P_{XY}}\!\left[\big{(}{k^{\mathcal{X}}}(X^{\prime},\cdot)-\mu_{P_{X}}\big{)}\otimes\big{(}{k^{\mathcal{Y}}}(Y^{\prime},\cdot)-\mu_{P_{Y}}\big{)}\right]\big{\rangle}_{\mathrm{HS}}$
=\ = $\mathbb{E}_{P_{XY},P_{XY}}\!\left[\big{\langle}\big{(}{k^{\mathcal{X}}}(X,\cdot)-\mu_{P_{X}}\big{)}\otimes\big{(}{k^{\mathcal{Y}}}(Y,\cdot)-\mu_{P_{Y}}\big{)},\big{(}{k^{\mathcal{X}}}(X^{\prime},\cdot)-\mu_{P_{X}}\big{)}\otimes\big{(}{k^{\mathcal{Y}}}(Y^{\prime},\cdot)-\mu_{P_{Y}}\big{)}\big{\rangle}_{\mathrm{HS}}\right]$
=\ = $\mathbb{E}_{P_{XY},P_{XY}}\!\left[\big{\langle}\big{(}{k^{\mathcal{X}}}(X,\cdot)-\mu_{P_{X}}\big{)},\big{(}{k^{\mathcal{X}}}(X^{\prime},\cdot)-\mu_{P_{X}}\big{)}\big{\rangle}_{{\mathcal{H}}_{k^{\mathcal{X}}}}\big{\langle}\big{(}{k^{\mathcal{Y}}}(Y,\cdot)-\mu_{P_{Y}}\big{)},\big{(}{k^{\mathcal{Y}}}(Y^{\prime},\cdot)-\mu_{P_{Y}}\big{)}\big{\rangle}_{{\mathcal{H}}_{k^{\mathcal{Y}}}}\right]$
=\ = $\mathbb{E}_{P_{XY},P_{XY}}\Big{[}\Big{(}{k^{\mathcal{X}}}(X,X^{\prime})-\mathbb{E}_{X}\!\left[{k^{\mathcal{X}}}(X,X^{\prime})\right]-\mathbb{E}_{X^{\prime}}\!\left[{k^{\mathcal{X}}}(X,X^{\prime})\right]+\mathbb{E}_{X,X^{\prime}}\!\left[{k^{\mathcal{X}}}(X,X^{\prime})\right]\Big{)}$
$\Big{(}{k^{\mathcal{Y}}}(Y,Y^{\prime})-\mathbb{E}_{Y}\!\left[{k^{\mathcal{Y}}}(Y,Y^{\prime})\right]-\mathbb{E}_{Y^{\prime}}\!\left[{k^{\mathcal{Y}}}(Y,Y^{\prime})\right]+\mathbb{E}_{Y,Y^{\prime}}\!\left[{k^{\mathcal{Y}}}(Y,Y^{\prime})\right]\Big{)}\Big{]}$
=\ = $\mathbb{E}_{P_{XY},P_{XY}}\Big{[}{k^{\mathcal{X}}}(X,X^{\prime})\Big{(}{k^{\mathcal{Y}}}(Y,Y^{\prime})-\mathbb{E}_{Y}\!\left[{k^{\mathcal{Y}}}(Y,Y^{\prime})\right]-\mathbb{E}_{Y^{\prime}}\!\left[{k^{\mathcal{Y}}}(Y,Y^{\prime})\right]+\mathbb{E}_{Y,Y^{\prime}}\!\left[{k^{\mathcal{Y}}}(Y,Y^{\prime})\right]\Big{)}\Big{]}$
=\ = $\mathbb{E}_{P_{XY},P_{XY}}\Big{[}{k^{\mathcal{X}}}(X,X^{\prime}){k^{\mathcal{Y}}}(Y,Y^{\prime})\Big{]}-2\,\mathbb{E}_{P_{XY}}\Big{[}\mathbb{E}_{P_{X}}[{k^{\mathcal{X}}}(X,X^{\prime})]\mathbb{E}_{P_{Y}}[{k^{\mathcal{Y}}}(Y,Y^{\prime})]\Big{]}$ (18)
@@
for kernels ${k^{\mathcal{X}}}$ and ${k^{\mathcal{Y}}}$ on ${\mathcal{X}}\times{\mathcal{X}}$ and ${\mathcal{Y}}\times{\mathcal{Y}}$, respectively, where we have used the property of the cross-covariance operator shown in Equation 2. We also mention the related conditional HSIC quantities of Zhang et al. (2011) and Pogodin et al. (2022, 2024).
@@
where $K^{\mathcal{X}}=\big{(}K^{\mathcal{X}}_{ij}\big{)}_{1\leq i,j\leq N}$, $K^{\mathcal{Y}}=\big{(}K^{\mathcal{Y}}_{ij}\big{)}_{1\leq i,j\leq N}$ and $H=I-\frac{1}{N}\bm{1}\bm{1}^{\top}$ is the centering matrix with $I$ the identity matrix and $\bm{1}$ a vector of ones, and where $\mathrm{tr}$ denotes the trace of a matrix. Finally, we present a one-sample second-order V-statistic for the HSIC, which does not consider all possible terms (unlike the one presented above) but which is useful to construct efficient estimators with faster computation time, as discussed in Section 3,
@@
For translation-invariant kernel $k_{\lambda}^{\mathcal{X}}$ and $k_{\mu}^{\mathcal{Y}}$ with bandwidths $\lambda$ and $\mu$, Equation 3 implies [^5] We use the convention that $(\cdot)_{\lambda,\mu}$ denotes $(\cdot)_{k_{\lambda}^{\mathcal{X}}\times k_{\mu}^{\mathcal{Y}}}$.
$\mathrm{HSIC}^{2}_{\lambda,\mu}\to 0\textrm{ when }\lambda\to 0\textrm{ or }\mu\to 0\textrm{ or }\lambda\to\infty\textrm{ or }\mu\to\infty,$ (28)
$V_{\mathrm{HSIC}^{2}_{\lambda,\mu}}\to\frac{1}{N}-\frac{1}{N^{2}}\textrm{ when }\lambda\to 0\textrm{ and }\mu\to 0,\ \ \textrm{ and }\ \ V_{\mathrm{HSIC}^{2}_{\lambda,\mu}}\to 0\textrm{ when }\lambda\to\infty\textrm{ or }\mu\to\infty,$
$U_{\mathrm{HSIC}^{2}_{\lambda,\mu}}\to 0\textrm{ when }\lambda\to 0\textrm{ or }\mu\to 0\textrm{ or }\lambda\to\infty\textrm{ or }\mu\to\infty.$
@@
=\ = $\mathbb{E}_{P_{XY},P_{XY}}[{k^{\mathcal{X}}}(X,X^{\prime}){k^{\mathcal{Y}}}(Y,Y^{\prime})]-2\mathbb{E}_{P_{XY},P_{X}P_{Y}}[{k^{\mathcal{X}}}(X,X^{\prime}){k^{\mathcal{Y}}}(Y,Y^{\prime})]+\mathbb{E}_{P_{X}P_{Y},P_{X}P_{Y}}[{k^{\mathcal{X}}}(X,X^{\prime}){k^{\mathcal{Y}}}(Y,Y^{\prime})]$
=\ = $\mathbb{E}_{P_{XY},P_{XY}}\Big{[}{k^{\mathcal{X}}}(X,X^{\prime}){k^{\mathcal{Y}}}(Y,Y^{\prime})\Big{]}-2\,\mathbb{E}_{P_{XY}}\Big{[}\mathbb{E}_{P_{X}}[{k^{\mathcal{X}}}(X,X^{\prime})]\mathbb{E}_{P_{Y}}[{k^{\mathcal{Y}}}(Y,Y^{\prime})]\Big{]}$
$+\mathbb{E}_{P_{X},P_{X}}\Big{[}{k^{\mathcal{X}}}(X,X^{\prime})\Big{]}\mathbb{E}_{P_{Y},P_{Y}}\Big{[}{k^{\mathcal{Y}}}(Y,Y^{\prime})\Big{]}$
=\ = $\mathrm{HSIC}^{2}_{{k^{\mathcal{X}}},{k^{\mathcal{Y}}}}(P_{XY}).$ (29)
@@
Now, consider the two-sample problem again with distributions $P,Q$, where for clarity we use variables $A,A^{\prime}\overset{\mathrm{i.i.d.}}{\sim}P$ and $B,B^{\prime}\sim Q$. Construct a joint distribution $P_{XY}$ with marginal $P_{X}=w_{P}P+w_{Q}Q$ a mixture of $P$ and $Q$ with positive weights $w_{P}+w_{Q}=1$, and set $Y=1$ if $X$ is drawn from $P$, or $Y=-1$ if $X$ is drawn from $Q$. For the labels, use the indicator kernel $k^{\mathcal{Y}}(y,y^{\prime})=\bm{1}(y=y^{\prime})$. For the data, we simply set the kernel to be the one used for two-sample testing, that is $k^{\mathcal{Y}}(x,x^{\prime})=k(x,x^{\prime})$. Then, we observe that
@@
Noting that $k_{\lambda}(\cdot,\cdot)\to\bm{1}(\cdot=\cdot)$ as the bandwidth $\lambda$ shrinks to 00, we also have
@@ 2.3 KSD: Kernel Stein Discrepancy
We present the KSD constructions of Chwialkowski et al. (2016) and Liu et al. (2016), more precisely, we follow the notation of the former. Let ${\mathcal{H}}$ be an RKHS in $\mathrm{Func}\!\left(\mathbb{R}^{d}\to\mathbb{R}\right)$ with reproducing kernel $k$. Denote by ${\mathcal{H}}^{d}$ the product RKHS consisting of elements of the form ${\bm{f}}=(f_{1},\dots,f_{d})$ with $f_{i}\in{\mathcal{H}}$ for $i=1,\dots,d$, with the associated inner product $\langle{\bm{f}},{\bm{g}}\rangle_{{\mathcal{H}^{d}}}=\sum_{i=1}^{d}\langle f_{i},g_{i}\rangle_{{\mathcal{H}}}$ for all ${\bm{f}},{\bm{g}}\in{\mathcal{H}^{d}}$. Note that elements of ${\mathcal{H}^{d}}$ can be seen as elements of $\mathrm{Func}\!\left(\mathbb{R}^{d}\to\mathbb{R}^{d}\right)$. The aim is to express the Stein operator with ${\mathcal{F}}\subseteq{\mathcal{H}^{d}}$ in terms of the kernel $k$, which will then give a closed-form expression to compute the Stein discrepancy. First, we recall the reproducing property of the kernel $k$ which implies that (Equation 1)
@@
The last equality holds under the Bochner integrability condition $\mathbb{E}_{Q}\!\left[\|{\bm{\xi}}_{P}(X)\|_{{\mathcal{H}^{d}}}\right]<\infty$ which itself holds provided by $\mathbb{E}_{Q}[h_{P}(X,X)]<\infty$ since $\mathbb{E}_{Q}\!\left[\|{\bm{\xi}}_{P}(X)\|_{{\mathcal{H}^{d}}}\right]\leq\sqrt{\mathbb{E}_{Q}\!\left[\|{\bm{\xi}}_{P}(X)\|_{{\mathcal{H}^{d}}}^{2}\right]}=\sqrt{\mathbb{E}_{Q}[h_{P}(X,X)]}$ as shown by Chwialkowski et al. (2016, Theorem 2.1). The Stein discrepancy with =${\mathcal{F}}=\{{\bm{f}}\in{\mathcal{H}^{d}}:\|{\bm{f}}\|_{{\mathcal{H}^{d}}}\leq 1\}$, which is referred to as KSD for Kernel Stein Discrepancy, is then equal to
@@
=$\overset{(\star)}{=}\mathbb{E}_{Q,Q}\!\left[\langle{\bm{\xi}}_{P}(X),{\bm{\xi}}_{P}(Y)\rangle_{{\mathcal{H}^{d}}}\right]=\mathbb{E}_{Q,Q}[h_{P}(X,Y)]$
@@
where $\langle\nabla k(x,\cdot),\nabla k(y,\cdot)\rangle_{{\mathcal{H}^{d}}}=\sum_{i=1}^{d}\big{\langle}\frac{\partial}{\partial x_{i}}k(x,\cdot),\frac{\partial}{\partial y_{i}}k(y,\cdot)\big{\rangle}_{{\mathcal{H}}}=\sum_{i=1}^{d}\frac{\partial^{2}}{\partial x_{i}\partial y_{i}}k(x,y)$ by the reproducing property with derivatives of Equation 1 (see also Steinwart and Christmann (2008, Lemma 4.34)). Again, the equation ()($\star$)( ⋆ ) holds under Bochner integrability which is satisfied when $\mathbb{E}_{Q}[h_{P}(X,X)]<\infty$. Stein’s identity gives $\mathbb{E}_{P}\!\big{[}(\mathcal{A}_{P}^{\mathcal{L}}{\bm{f}})(x)\big{]}=0$ for all ${\bm{f}}\in{\mathcal{F}}$, which implies that $\mathbb{E}_{P}[{\bm{\xi}}_{P}(X)]=0$, and hence $\mathbb{E}_{P}[h_{P}(X,\cdot)]=0$.
Writing ${\bm{\delta}}(x)={\bm{s}_{P}}(x)-{\bm{s}_{Q}}(x)$ for the difference in score, we recall that Equation 38 gives that for any ${\bm{f}}\in{\mathcal{F}}$ we have
@@
where the first and second equalities hold by Equation 44 with ${\bm{f}}_{1}(X)=k(X,Y){\bm{\delta}}(Y)$ for fixed $Y$ and ${\bm{f}}_{2}(Y)=k(X,Y){\bm{s}_{P}}(X)+\nabla_{\!X}k(X,Y)$ for fixed $X$, respectively, giving
@@
$(\mathcal{A}_{P}^{\mathcal{L}}{\bm{f}}_{2})(Y)=$ $\Big{(}k(X,Y){\bm{s}_{P}}(X)+\nabla_{\!X}k(X,Y)\Big{)}^{\top}{\bm{s}_{P}}(Y)+\nabla_{\!Y}^{\top}\Big{(}k(X,Y){\bm{s}_{P}}(X)+\nabla_{\!X}k(X,Y)\Big{)}$
=\ = $k(X,Y){\bm{s}_{P}}(X)^{\top}{\bm{s}_{P}}(Y)+\big{(}\nabla_{\!X}k(X,Y)\big{)}^{\top}{\bm{s}_{P}}(Y)+\big{(}\nabla_{\!Y}k(X,Y)\big{)}^{\top}{\bm{s}_{P}}(X)+\nabla_{\!Y}^{\top}\big{(}\nabla_{\!X}k(X,Y)\big{)}$
=\ = $h_{P}(X,Y)$

### v2

@@ Abstract
This paper corresponds to the introduction of my PhD thesis (Schrab, 2025a, Chapter 2) and is presented as a standalone article to introduce the reader to kernel discrepancies estimators. First, in Section˜1, we define kernels, Reproducing Kernel Hilbert Spaces, mean embeddings and cross-covariance operators, and present kernel properties such as characteristicity, universality and translation invariance. Then, in Section˜2, we introduce the Maximum Mean Discprecancy, the Hilbert–Schmidt Independence Criterion, and the Kernel Stein Discrepancy, as well as their estimators, and we discuss the importance of the choice of kernel for such measures. We then introduce a collection of statistics in Section˜3, including the commonly-used complete statistics, as well as their incomplete counterparts which trade accuracy for computational efficiency. Finally, in Section˜4, we construct adaptive estimators combining multiple statistics with various kernels, which is one method to address the problem of kernel selection.
@@ 1 Kernels
  • $1\Longrightarrow 2]$ Every kernel is a positive definite kernel as $\sum_{i=1}^{n}\sum_{j=1}^{n}c_{i}c_{j}k(x_{i},x_{j})=\|\sum_{i=1}^{n}c_{i}\phi(x_{i})\|_{{\mathcal{H}}}\geq 0$.
  • $2\Longrightarrow 1]$ Every positive definite kernel is guaranteed to be an inner product between features in an Hilbert space (Steinwart and Christmann, 2008, Theorem 4.16).
  • $2\Longrightarrow 3]$ Given a positive definite kernel $k$ on ${\mathcal{X}}\times{\mathcal{X}}$, there exists a unique Hilbert space ${\mathcal{H}}_{k}$ of real-valued functions on $\mathcal{X}$} for which $k$ is the reproducing kernel, ${\mathcal{H}}_{k}$ is the (unique) RKHS associated to $k$ (Aronszajn, 1950, Moore-Aronsajn Theorem).
  • $3\Longrightarrow 1]$ A reproducing kernel is a kernel with feature map $\phi(x)=k(x,\cdot)$ as $\langle k(x,\cdot),k(y,\cdot)\rangle_{{\mathcal{H}}}=k(x,y)$.
@@
  • $3\Longrightarrow 4]$ Using the reproducing property followed by Cauchy–Schwartz inequality, we obtain $|f(x)|\leq\sqrt{k(x,x)}\|f\|_{{\mathcal{H}}}$ for all $f\in{\mathcal{H}}$ and all $x\in{\mathcal{X}}$.
  • $4\Longrightarrow 3]$ By Riesz Representation Theorem (Rudin, 1987, Theorem 4.12), for all $x\in{\mathcal{X}}$ there exists $g_{x}\in{\mathcal{H}}$ such that $f(x)=\langle f,g_{x}\rangle_{{\mathcal{H}}}$ for all $f\in{\mathcal{H}}$. The function $k(x,y)\coloneqq g_{x}(y)$ satisfies the two properties of definition 3 and hence is a repoducing kernel.
@@
[…]
@@ 2 Kernel discrepancies
We introduce the Maximum Mean Discrepancy (MMD) in Section˜2.1, the Hilbert–Schmidt Independence Criterion (HSIC) in Section˜2.2, and the Kernel Stein Discrepancy (KSD) in Section˜2.3.
@@ 2.1 MMD: Maximum Mean Discrepancy
[…]
@@
for $x,x^{\prime},y,y^{\prime}\in\mathbb{R}^{d}$. Writing the estimator $V_{\mathrm{MMD}^{2}_{k}}$ as a two-sample V-statistic can be theoretically appealing but we stress that it can in fact be computed in quadratic time using Equation˜7. The V-statistic incorporates the terms $\{k(X_{i},Y_{i}):i=1,\dots,n\}$ and, hence, is biased. We also point out that the V-statistic is always non-negative, and taking its square root yields an estimator of the (non-squared) MMD. The global sensitivity of this MMD statistic is studied in Kim and Schrab (2023, Lemma 5), which provides robustness guarantees (Schrab and Kim, 2025).
@@
An unbiased estimator of the squared MMD (Gretton et al., 2012, Lemma 6) naturally arises from Equation˜6 as
@@
The MMD U-statistic expression of Equation˜12 cannot be expressed as a single vector-matrix-vector product, instead it needs to be computed as
@@
$\widetilde{U}_{\mathrm{MMD}^{2}_{k}}=\frac{1}{n(n-1)}v^{\top}\bar{K}^{ZZ}v$ (16)
where $\bar{K}^{ZZ}$ is the matrix $K^{ZZ}$ as defined in Equation˜10 but with diagonal entries set to zero, and $v\coloneqq(\bm{1}_{n},-\bm{1}_{n})$ is a vector of signed ones. As a result of the U-statistic being unbiased, it is not always non-negative and hence cannot be used to estimate the MMD by taking its square root (as for the V-statistic).
@@
When using the translation-invariant kernel $k_{\lambda}$, Equation˜3 implies [^3] We use the convention that $(\cdot)_{\lambda}$ denotes $(\cdot)_{k_{\lambda}}$ throughout the thesis.
@@ 2.2 HSIC: Hilbert–Schmidt Independence Criterion
[…]
@@
for kernels $k^{\mathcal{X}}$} and $k^{\mathcal{Y}}$} on ${\mathcal{X}}\times{\mathcal{X}}$ and ${\mathcal{Y}}\times{\mathcal{Y}}$, respectively, where we have used the property of the cross-covariance operator shown in Equation˜2. We also mention the related conditional HSIC quantities of Zhang et al. (2011) and Pogodin et al. (2022, 2024).
@@
where $K^{\mathcal{X}}=\big{(}K^{\mathcal{X}}_{ij}\big{)}_{1\leq i,j\leq N}$, $K^{\mathcal{Y}}=\big{(}K^{\mathcal{Y}}_{ij}\big{)}_{1\leq i,j\leq N}$ and $H=I-\frac{1}{N}\bm{1}\bm{1}^{\top}$ is the centering matrix with $I$ the identity matrix and $\bm{1}$ a vector of ones, and where $\mathrm{tr}$ denotes the trace of a matrix. Finally, we present a one-sample second-order V-statistic for the HSIC, which does not consider all possible terms (unlike the one presented above) but which is useful to construct efficient estimators with faster computation time, as discussed in Section˜3,
@@
For translation-invariant kernel $k_{\lambda}^{\mathcal{X}}$ and $k_{\mu}^{\mathcal{Y}}$ with bandwidths $\lambda$ and $\mu$, Equation˜3 implies [^5] We use the convention that $(\cdot)_{\lambda,\mu}$ denotes $(\cdot)_{k_{\lambda}^{\mathcal{X}}\times k_{\mu}^{\mathcal{Y}}}$.
[…]
@@
[…]
@@
Now, consider the two-sample problem again with distributions $P,Q$, where for clarity we use variables $A,A^{\prime}\overset{\mathrm{i.i.d.}}{\sim}P$ and $B,B^{\prime}\sim Q$. Construct a joint distribution $P_{XY}$ with marginal $P_{X}=w_{P}P+w_{Q}Q$ a mixture of $P$ and $Q$ with positive weights $w_{P}+w_{Q}=1$, and set $Y=1$ if $X$ is drawn from $P$, or Y=1Y=-1 if $X$ is drawn from $Q$. For the labels, use the indicator kernel $k^{\mathcal{Y}}(y,y^{\prime})=\bm{1}(y=y^{\prime})$. For the data, we simply set the kernel to be the one used for two-sample testing, that is $k^{\mathcal{Y}}(x,x^{\prime})=k(x,x^{\prime})$. Then, we observe that
@@
Noting that $k_{\lambda}(\cdot,\cdot)\to\bm{1}(\cdot=\cdot)$ as the bandwidth $\lambda$ shrinks to 0, we also have
@@ 2.3 KSD: Kernel Stein Discrepancy
We present the KSD constructions of Chwialkowski et al. (2016) and Liu et al. (2016), more precisely, we follow the notation of the former. Let $\mathcal{H}$} be an RKHS in $\mathrm{Func}\!\left(\mathbb{R}^{d}\to\mathbb{R}\right)$ with reproducing kernel $k$. Denote by ${\mathcal{H}}^{d}$ the product RKHS consisting of elements of the form ${\bm{f}}=(f_{1},\dots,f_{d})$ with $f_{i}\in{\mathcal{H}}$ for $i=1,\dots,d$, with the associated inner product $\langle{\bm{f}},{\bm{g}}\rangle_{{\mathcal{H}^{d}}}=\sum_{i=1}^{d}\langle f_{i},g_{i}\rangle_{{\mathcal{H}}}$ for all ${\bm{f}},{\bm{g}}\in{\mathcal{H}^{d}}$. Note that elements of $\mathcal{H}^{d}$} can be seen as elements of $\mathrm{Func}\!\left(\mathbb{R}^{d}\to\mathbb{R}^{d}\right)$. The aim is to express the Stein operator with ${\mathcal{F}}\subseteq{\mathcal{H}^{d}}$ in terms of the kernel $k$, which will then give a closed-form expression to compute the Stein discrepancy. First, we recall the reproducing property of the kernel $k$ which implies that (Equation˜1)
@@
The last equality holds under the Bochner integrability condition $\mathbb{E}_{Q}\!\left[\|{\bm{\xi}}_{P}(X)\|_{{\mathcal{H}^{d}}}\right]<\infty$ which itself holds provided by $\mathbb{E}_{Q}[h_{P}(X,X)]<\infty$ since $\mathbb{E}_{Q}\!\left[\|{\bm{\xi}}_{P}(X)\|_{{\mathcal{H}^{d}}}\right]\leq\sqrt{\mathbb{E}_{Q}\!\left[\|{\bm{\xi}}_{P}(X)\|_{{\mathcal{H}^{d}}}^{2}\right]}=\sqrt{\mathbb{E}_{Q}[h_{P}(X,X)]}$ as shown by Chwialkowski et al. (2016, Theorem 2.1). The Stein discrepancy with ${\mathcal{F}}=\{{\bm{f}}\in{\mathcal{H}^{d}}:\|{\bm{f}}\|_{{\mathcal{H}^{d}}}\leq 1\}$, which is referred to as KSD for Kernel Stein Discrepancy, is then equal to
@@
$\overset{(\star)}{=}\mathbb{E}_{Q,Q}\!\left[\langle{\bm{\xi}}_{P}(X),{\bm{\xi}}_{P}(Y)\rangle_{{\mathcal{H}^{d}}}\right]=\mathbb{E}_{Q,Q}[h_{P}(X,Y)]$
@@
where $\langle\nabla k(x,\cdot),\nabla k(y,\cdot)\rangle_{{\mathcal{H}^{d}}}=\sum_{i=1}^{d}\big{\langle}\frac{\partial}{\partial x_{i}}k(x,\cdot),\frac{\partial}{\partial y_{i}}k(y,\cdot)\big{\rangle}_{{\mathcal{H}}}=\sum_{i=1}^{d}\frac{\partial^{2}}{\partial x_{i}\partial y_{i}}k(x,y)$ by the reproducing property with derivatives of Equation˜1 (see also Steinwart and Christmann (2008, Lemma 4.34)). Again, the equation $(\star)$ holds under Bochner integrability which is satisfied when $\mathbb{E}_{Q}[h_{P}(X,X)]<\infty$. Stein’s identity gives $\mathbb{E}_{P}\!\big{[}(\mathcal{A}_{P}^{\mathcal{L}}{\bm{f}})(x)\big{]}=0$ for all $\bm{f}$$\in{\mathcal{F}}$, which implies that $\mathbb{E}_{P}[{\bm{\xi}}_{P}(X)]=0$, and hence $\mathbb{E}_{P}[h_{P}(X,\cdot)]=0$.
Writing ${\bm{\delta}}(x)={\bm{s}_{P}}(x)-{\bm{s}_{Q}}(x)$ for the difference in score, we recall that Equation˜38 gives that for any $\bm{f}$$\in{\mathcal{F}}$ we have
@@
where the first and second equalities hold by Equation˜44 with ${\bm{f}}_{1}(X)=k(X,Y){\bm{\delta}}(Y)$ for fixed $Y$ and ${\bm{f}}_{2}(Y)=k(X,Y){\bm{s}_{P}}(X)+\nabla_{\!X}k(X,Y)$ for fixed $X$, respectively, giving
@@
$(\mathcal{A}_{P}^{\mathcal{L}}{\bm{f}}_{2})(Y)=\$$\Big{(}k(X,Y){\bm{s}_{P}}(X)+\nabla_{\!X}k(X,Y)\Big{)}^{\top}{\bm{s}_{P}}(Y)+\nabla_{\!Y}^{\top}\Big{(}k(X,Y){\bm{s}_{P}}(X)+\nabla_{\!X}k(X,Y)\Big{)}$
[…]